If you want to make 200.00 mL of a solution that is 1.75 ppm and you have a 12.0 ppm stock solution, how much of the stock solution should you measure out and dilute with water?
Answer #1

Answer :

Concentration of stock solution ( C1) = 12 ppm

Volume of stock solution (V1) = ?

Concentration of given prepared solution ( C2) = 1.75 ppm

Volume of given prepared solution ( V2) = 200 ml

From Molarity equation :

C1 * V1 = C2 * V2

V1 = C2 * V2 / C1 = 1.75 * 200 / 12 = 17.5 ml

Volume of stock solution required = 17.5 ml

Volume of water required = volume of given prepared solution - volume of stock solution = 200 - 17.5 = 182.5 ml

Volume of water required = 182.5 ml